Given ?2 + ? + ?2 = 9 Find an equation of the tangent plane and a set of parametric equations of the normal line at the point P(1, 4, 2).
Answer #1

x²+4 + Z² = 9 P(1,4,2) 2 let f(2.9,7)= x2 + y + z2 , / If = {2x, 1 27) IF(1,4,2)= < 2 1,4> tangent plane at p has equation 2(